Solve each of the following.

7. Solve each of the following: (a) y - 3y + 2y = 0 (d) y - 4y + 5y = 0 (b) y - 10y + 25y = 0. (C) y + 3y - 5y = 0 (e
Answer #1

a -34tayao consider auxillon egon m² - 3 mta-o m²-am-mtazo mm-2)-ilm-2)=0 (m-)) (m-2) 20 → mH com- do Malf m=2 noots are dm² +3m-5=0 - - 3 + vg+20 m = -3 + 129 m 2 mis - 3#vag me -3-12 & YM) - ciemint ca e man - 3+ 129 x ya) = le -3 -128 x & le 2@ yllty=0 auxillon eon m²+1=0 m = 19 a=096=1 roots are complex y(x) = x(cleosa+(asina) 4x) = cicosat (a sinn I 7 0 1 cicos Itmuliply oy b & 6 by the then add (2 + 2) 2 = 2 2 2 s (2=2 2 hm 6 cita = -(2 82 il- – 43 SV2 V 2 - U Yin) - - 43 coox of 4 12